# Channel Choice and Customer Value

### v1

@@ Appendix A Data cleaning
Table A1: Item categories with substantially different availability time via online and offline channels (YY-MM-DD)
@@
Appendix B Appendix to Descriptive Analysis
In this section, we provide some descriptive analyses summarizing the relationship between channel choice and consumers’ purchase behaviors. In particular, we are interested in quantifying whether the value/revenue generated by a customer is systematically correlated with their channel choice. In $\S$§B.1, we first present some cross-sectional evidence linking channel choice to customer value. Next, in $\S$§B.2, we present a within-user analysis that can control for consumer-specific variables.
B.1 Cross-sectional Comparison
Table A2: Frequency distribution of gender – offline-only/online-only/multichannel
Gender Online-only (%) Offline-only (%) Multichannel (%) Overall (%)
No Information 41.95 37.24 24.50 34.75
Female 37.75 35.23 49.08 39.62
Male 20.30 27.53 26.42 25.63
Total 100 100 100 100
Figure A1: Cumulative distribution function (CDF) of Age/Monthly Household Income – offline-only, online-only,multichannel
We start by categorizing consumers into three mutually exclusive groups:
  • $\bullet$
    Multichannel consumers – consumers who have purchased from both offline and online channels at least once during the observation period.
  • $\bullet$
    Offline-only consumers – consumers who have exclusively purchased from physical stores during the observation period.
  • $\bullet$
    Online-only consumers – consumers who have only purchased using the online channels (website and/or mobile app) during the observation period.
In total, our data consists of 1,642,859 (27.7%) multichannel consumers, 2,986,541 (50.3%) offline-only consumers, and 1,304,848 (22.0%) online-only consumers.
Table A3: Customer-Month Summary Statistics By Customer Types – Multichannel/Offline-only/Online-only
Variables Mean Std Min 25% 50% 75% Max Count
Multichannel Customers
Spend 398.16 805.24 0 0 0 548.07 301,134.12 56,151,374
Quantity 3.15 9.69 0 0 0 3.00 4,301.00 56,151,374
Order 0.76 1.18 0 0 0 1.00 236.00 56,151,374
UniqueItem 1.74 3.20 0 0 0 2.00 355.00 56,151,374
UniqueBrand 1.39 2.32 0 0 0 2.00 127.00 56,151,374
UniqueSubcategory 1.20 1.87 0 0 0 2.00 43.00 56,151,374
UniqueCategory 1.01 1.46 0 0 0 2.00 16.00 56,151,374
Offline-only Customers
Spend 221.46 738.65 0 0 0 226.34 2,340,617.16 56,259,534
Quantity 1.96 11.38 0 0 0 2.00 44,158.00 56,259,534
Order 0.50 0.91 0 0 0 1.00 430.00 56,259,534
UniqueItem 1.25 2.82 0 0 0 2.00 4,191.00 56,259,534
UniqueBrand 1.01 1.96 0 0 0 1.00 332.00 56,259,534
UniqueSubcategory 0.89 1.62 0 0 0 1.00 64.00 56,259,534
UniqueCategory 0.76 1.29 0 0 0 1.00 18.00 56,259,534
Online-only Customers
Spend 236.49 539.39 0 0 0 311.22 110,682.95 11,337,544
Quantity 1.33 6.21 0 0 0 1.00 4,800.00 11,337,544
Order 0.42 0.68 0 0 0 1.00 130.00 11,337,544
UniqueItem 0.72 1.49 0 0 0 1.00 141.00 11,337,544
UniqueBrand 0.62 1.15 0 0 0 1.00 68.00 11,337,544
UniqueSubcategory 0.56 0.98 0 0 0 1.00 25.00 11,337,544
UniqueCategory 0.51 0.84 0 0 0 1.00 11.00 11,337,544
We find that these three groups are systematically different from each other in terms of demographics (see Table A2 and Figure A1). Multichannel consumers are more likely to be female compared to offline and online consumers. Online-only consumers tend to be younger than both offline and multichannel consumers. Further, offline-only consumers tend to be lower income while multichannel consumers tend to be the wealthiest. There are also significant differences in channel choice across geographic locations (see Figure A2). Overall, these differences suggest that there is significant self-selection in channel choice based on consumer attributes.
Figure A2: Fraction of multichannel customers to total customers within states
Next, we summarize the purchase behaviors of each of these groups in Table A3. The means shown in this table are calculated for each group $g$ (where $g$ is either multichannel, offline-only, or online-only) as follows: $\frac{\sum_{i=1}^{N_{g}}\sum_{t=1}^{T_{i}}X_{it}}{\sum_{i=1}^{N_{g}}T_{i}}$, where $i$ denotes a customer belonging to group $g$, $t$ denotes the year-month index, $T_{i}$ is the panel length of customer $i$, and $N_{g}$ is the total number of consumers in group $g$. As we can see, multichannel customers tend to spend more, and buy higher quantities and larger varieties of products compared to offline-only or online-only consumers. If we compare consumers who exclusively buy offline with those who exclusively buy online, we see that online-only consumers tend to spend more but purchase lower quantities and lower variety of items. This is mainly because online-only consumers tend to buy larger pack sizes of pet foods and more expensive products compared to offline-only consumers.
However, these patterns could be explained by the differences in observed consumer attributes across the three channels. To examine if this is the case, we estimate the following linear regression:
$y_{it}=CustomerType_{i}+Gender_{i}+Age_{i}+HouseholdIncome_{i}+Zipcode_{i}+YearMonth_{t}+\epsilon_{it}$ (A1)
where $y_{it}$ can denote one of the purchase metrics (e.g. Spend, Orders, UniqueCategories, etc.) of customer $i$ in month $t$. The main explanatory variable of interest is $CustomerType_{i}$, which is a dummy variable that denotes the customer’s channel choice: multichannel, offline-only, or online-only. We specify online-only as the baseline category. In addition, we include other control variables related to customer demographics (Gender, Age, Zipcode, and HouseholdIncome) for the reasons discussed earlier, i.e., to control for individual characteristics that can affect purchase behavior. Finally, we also control for time-varying unobserved shocks in period $t$ that can affect purchase behavior (such as seasonality and economic factors) using a period fixed effect ($YearMonth_{t}$).
Table A4: Cross-sectional Comparison of Customers’ Monthly Spend
DV: Spend
Model: (1) (2)
Variables
Constant 236.5 (0.412)
[0.000]
Multichannel 161.7 (0.595) 149.0 (0.610)
[0.000] [0.000]
Offline-Only -15.0 (0.509) -31.2 (0.561)
[0.000] [0.000]
Control Variables
Age Yes
Gender Yes
Zipcode Yes
Household Income Yes
Fixed-effects
YearMonth Yes
Fit statistics
Observations 123,748,697 123,748,697
R2 0.01307 0.04867
Within R2 0.01165
Clustered (Customer) standard‐errors are presented in standard brackets,
while p‑values are shown in square brackets.
Table A4 presents the difference in total spend of the multichannel and offline-only customers relative to the baseline category (i.e., online-only customers). In column (1), we use only customer type as the explanatory variable and find that spending patterns across customer types are the same as those discussed above. Next, in column (2), we estimate the same model with all the controls/fixed effects and find that the effects persist even after controlling for selection using the observed seasonality and demographic variables. Specifically, we find that, compared to offline-only consumers, multichannel consumers spend about 180.2 MCU more per month. Interestingly, the magnitude of this effect is almost the same in both models (1) and (2), i.e., the addition of controls and fixed effects did not materially change this result. Further, we replicate the specification in column (2) of Table A4 with other purchase metrics in Table A5 and find that the results are consistent with the observations from Table A3.
Table A5: Cross-sectional Comparisons of Multichannel and Offline-only relative to Online-only customers
DVs: Quantities Orders Unique Items Unique Brands Unique Subcategories Unique Categories
Model: (1) (2) (3) (4) (5) (6)
Variables
Multichannel 1.57 (0.009) 0.294 (0.001) 0.901 (0.002) 0.686 (0.001) 0.564 (0.001) 0.439 (0.001)
[0.000] [0.000] [0.000] [0.000] [0.000] [0.000]
Offline-Only 0.260 (0.008) 0.006 (0.001) 0.373 (0.002) 0.277 (0.001) 0.225 (0.001) 0.162 (0.001)
[0.000] [0.000] [0.000] [0.000] [0.000] [0.000]
Fixed-effects
YearMonth Yes Yes Yes Yes Yes Yes
Age Yes Yes Yes Yes Yes Yes
Gender Yes Yes Yes Yes Yes Yes
Zip code Yes Yes Yes Yes Yes Yes
Household Income Yes Yes Yes Yes Yes Yes
Fit statistics
Observations 123,748,697 123,748,697 123,748,697 123,748,697 123,748,697 123,748,697
R2 0.02914 0.06216 0.04918 0.05339 0.05514 0.05695
Within R2 0.00382 0.01715 0.01004 0.01162 0.01193 0.01229
These findings align with previous literature such as Kumar and Venkatesan (2005); Thomas and Sullivan (2005); Kushwaha and Shankar (2013), who find that consumers who use multiple channels are more valuable than consumers who only purchase via one channel (though the set of channels studied in that literature are different). Given that offline-only consumers are the lowest spenders and multichannel consumers are the heaviest spenders, a natural question for the firm is whether it can increase revenues and profits by nudging offline-only consumers to also sign up for its online channels, or if more consumers sign up for online channels due to economic or environmental changes. However, the results from Table A4 could be a combination of the effect of having access to multiple channels as well as unobserved but systematic differences across consumers who use the two channels, and the reasons why they chose to adopt the online channels. For example, users who are loyal to the firm and/or heavy consumers are more likely to use both offline and online channels. As such, converting an offline-only consumer to use both offline and online channels may not necessarily lead to the same incremental gains. Further, a cross-sectional comparison doesn’t inform us about the changes in purchase behavior that occur after a consumer starts using the online channels in addition to offline.
B.2 Within-consumer Comparison
Table A6: Customers’ Spend Before/After Adopting the Online Channel
DV: Spend
Model: (1) (2)
Variables
Constant 318.4 (0.621)
[0.000]
Post_online_adoption 147.4 (0.764) 124.2 (0.782)
[0.000] [0.000]
Fixed-effects
Customer Yes
YearMonth Yes
Fit statistics
Observations 34,226,864 34,226,864
R2 0.00754 0.34036
Within R2 0.00275
Clustered (Customer) standard‐errors are presented in standard brackets,
while p‑values are shown in square brackets.
Table A7: Customers’ Purchase Behaviors Before/After Adopting Online Shopping
DVs: Quantities Orders Unique Items Unique Brands Unique Subcategories Unique Categories
Model: (1) (2) (3) (4) (5) (6)
Variables
Post_online_adoption 1.05 (0.011) 0.289 (0.001) 0.279 (0.003) 0.214 (0.002) 0.179 (0.002) 0.168 (0.001)
[0.000] [0.000] [0.000] [0.000] [0.000] [0.000]
Fixed-effects
Customer Yes Yes Yes Yes Yes Yes
YearMonth Yes Yes Yes Yes Yes Yes
Fit statistics
Observations 34,226,864 34,226,864 34,226,864 34,226,864 34,226,864 34,226,864
R2 0.41355 0.32593 0.31713 0.29285 0.27776 0.27099
Within R2 0.00141 0.00663 0.00078 0.00085 0.00091 0.00131
Given the shortcomings of the cross-sectional analysis discussed above, we now perform a within-consumer comparison that allows us to control for some of the selection issues, and also study the changes in consumer purchase behavior after online shopping adoption. We focus on consumers who were offline-only at the beginning of our observation period and adopted the online channel sometime during the observation period. 898,413 consumers satisfy this criterion, and we use their data for our analysis. To cleanly identify the difference in spending before and after the adoption of online shopping, we exclude the observations of the month a customer decides to adopt online shopping. We then run the following regression on these consumers’ monthly spending:
$y_{it}=Customer_{i}+\beta_{0}Post\_online\_adoption_{it}+YearMonth_{t}+\epsilon_{it},$ (A2)
where $y_{it}$ refers to an outcome variables of interest, $Customer_{i}$ is a consumer fixed-effect,
$Post\_online\_adoption_{it}$ is an indicator for whether consumer $i$ had adopted online shopping before period $t$, and $YearMonth_{t}$ is a time period fixed effect, as before. Table A6 presents the results from this regression for Spend as the outcome variable. Model (1) shows the results without the customer and year-month fixed effects and model (2) shows the results with the fixed effects. We see that, on average, customers spend 124.2 MCU more after they adopt online shopping. Note that this result is significantly smaller than the 180.2 MCU from the previous cross-sectional analysis. These results are consistent with earlier work (studying different channel transitions) that shows that consumers who transition to shopping on multiple channels from previously using only a single channel tend to spend more (Venkatesan, Kumar, and Ravishanker 2007; Ansari, Mela, and Neslin 2008) post multichannel adoption.
We also perform a similar analysis on other metrics of interest (Quantity, No. of orders, Unique items, Brands, Categories, and Subcategories) and find that all of them show a positive increase after the adoption of online shopping (see Table A7). However, as with spend, the magnitude of these effects is much smaller than what we saw in the cross-sectional analysis.
However, while the within-consumer analysis allows us to study changes in behavior for the same consumers, there is still a strong assumption that the changes we observe between the before and after periods are entirely attributable to the adoption of online shopping, and there are no other time varying factors in play causing some of the changes in behavior we observe. As such, without including an appropriate control group of non-adopters, the effect size here could still be an inflated measure of the incremental value from converting an offline-only consumer to a multichannel consumer who also uses online channels. Moreover, these analyses do not shed light on whether the different reasons for the adoption of online shopping have any significant impact on a consumer’s purchase behavior in the post-adoption period.
Appendix C Appendix to Summary Statistics
@@
To characterize the pre-online-adoption purchase behavior of each group of customers, we compute several consumer-level metrics, such as AvgSpendPerMonth, AvgOrdersPerMonth, AvgUniqueCategoriesPerMonth, using data from the pre-adoption period specific to each case study. We summarize these variables for each group – event adopters, organic adopters, and offline-only – by calculating the group mean as follows: $\frac{1}{N_{g}}\sum_{i=1}^{N_{g}}(\frac{1}{T_{i}^{pre}}\sum_{t=1}^{T_{i}^{pre}}X_{it})$, where $i$ denotes a customer belonging to group $g$, $t$ denotes the year-month index, $T_{i}^{pre}$ denotes the length of the consumer $i$’s panel during the pre-adoption period, and $N_{g}$ is the number of customers in group $g$. We next summarize the post-online adoption behaviors in a similar way using the post-online adoption data under each case study. The post-online adoption means for each group $g$event adopters, organic adopters, and offline-only – are calculated as follows: $\frac{1}{N_{g}}\sum_{i=1}^{N_{g}}(\frac{1}{T_{i}^{post}}\sum_{t=1}^{T_{i}^{post}}X_{it})$, where $i$ denotes a customer belonging to group $g$, $t$ denotes the year-month index, $T_{i}^{post}$ denotes the length of the consumer $i$’s panel during the post-adoption period, and $N_{g}$ is the number of customers in group $g$.
C.1 Event 1: COVID-19
In this section, we present the descriptive statistics for customers in the sample for the COVID-19. We first summarize demographics in Table A8. Next, in Tables A9 and A10, we show customer-level summary statistics of average purchase variables over the pre-adoption and post-adoption period, respectively. Finally, in Table A11, we present more detailed summary statistics (e.g., median and quartiles) for all three groups during the post-online-adoption periods (i.e., between May 1st, 2020 and April 30, 2023). Note that the counts of organic adopters and covid adopters from Table A11 slightly differ from the raw count based on our cohorts’ definition because 469 (3.66% of) organic adopters and 933 (2.81% of) covid adopters churned before 2020-05-01, and therefore are not observed in the post-COVID-19 panel. Including these customers does not affect our identification of post-online-adoption ATT, and the estimation results remain the same.
@@ C.1 Event 1: COVID-19
C.2 Event 2: Black Friday
In this section, we present the descriptive statistics for customers in the sample for the Black Friday event. We first summarize demographics in Table A12. Next, in Tables A13 and A14, we show customer-level summary statistics of average purchase variables over the pre-adoption and post-adoption period, respectively. Finally, in Table A15, we present more detailed summary statistics (e.g., median and quartiles) for all three groups during the post-online-adoption periods (i.e., between December 1st, 2019 and October 31, 2020). Note that the counts of organic adopters and Black Friday adopters from Table A15 slightly differ from the raw count based on our cohorts’ definition because 79 (2.58% of) organic adopters and 24 (1.74% of) Black Friday adopters churned before 2019-12-01, and therefore are not observed in the post-Black Friday panel. Including these customers does not affect our identification of post-online-adoption ATT, and the estimation results remain the same.
@@ C.2 Event 2: Black Friday
C.3 Event 3: Loyalty Program
In this section, we present the descriptive statistics for customers in the sample for the loyalty program analysis. We first summarize demographics in Table A16. Next, in Tables A17 and A18, we show customer-level summary statistics of average purchase variables over the pre-adoption and post-adoption period, respectively. Finally, in Table A19, we present more detailed summary statistics (e.g., median and quartiles) for all three groups during the post-online-adoption periods (i.e., between October 1st, 2023 and March 31, 2024)
@@ C.3 Event 3: Loyalty Program
C.4 Loyalty Program Control Variables
@@
    %$\%Month\_WithActivation_{it}$: The proportion of months in which customer $i$ has activated the offer up to (but not including) month $t$, calculated as the total number of activation months up to $t$ divided by the number of months since first activation (excluding $t$ itself). For the first month of activation, this variable is defined as 0 (since the denominator is 0); all first-month effects are captured by $isFirstActivation_{it}$. Starting from the second month post-activation, this variable reflects the historical activation rate prior to month $t$.
Appendix D Appendix to the Profitability Analysis: Missing Value Imputation
@@
Appendix E Appendix to Mechanisms, Theoretical Underpinnings and Managerial Implications
E.1 Appendix to the Share of Offline-Online Spend
@@
Table A23: Differential Trends in the Share of Offline Spend (Estimation Results for Equation A5) – Event Adopters vs. Organic Adopters
@@
Table A20 presents the DID results of offline spend between adopters and offline-only customers. Table A21 and A22 shows the DID results of offline spend and share of offline spend between event adopters and organic adopters. Table A23 presents the estimation results for the linear trend model by comparing a type of event adopter with organic adopters. Table A24 presents the changes in share of offline spend by different age groups, showing senior adopters tend to have a slower rate of shifting their spend to the online channel.
In §7.1, we find that the share of offline spending is persistently higher for covid adopters compared to organic adopters in the post-adoption period, and the downward trend is steeper for organic adopters, indicating that covid adopters shift their spending to online channels slower than organic adopters. To formally test whether these differential trends in share of offline spend are systematic, we estimate the following model with linear time trend on the post-adoption monthly panel data for both groups:
@@
(a) Senior Adopters (Age 41$\geq$ 41≥ 41)
(b) Young Adopters (Age <41<41< 41)
E.2 Appendix to the Other Metrics
Finally, to further explore the differences in the purchase behavior of different groups of adopters relative to organic adopters post-online-adoption, we estimate the same DID model (as shown in Equation (3) and (4)) on other metrics of interest (Quantity, Number of orders, Unique items, Brands, Categories, and Subcategories). For each study (Covid, Black Friday, and Loyalty Program), we present results for both total and offline metrics. Specifically, total metrics are reported in Tables A25, A27, and A29, while offline metrics appear in Tables A26, A28, and A30, respectively.

### v2

@@ Appendix A Data cleaning
Table A1: Item categories with substantially different availability time via online and offline channels ($Y$-$M$-$D$)
@@
Appendix B Appendix to Summary Statistics
@@
To characterize the pre-online-adoption purchase behavior of each group of customers, we compute several consumer-level metrics, such as AvgSpendPerMonth, AvgOrdersPerMonth, AvgUniqueCategoriesPerMonth, using data from the pre-adoption period specific to each case study. We summarize these variables for each group – event adopters, organic adopters, and offline-only – by calculating the group mean as follows: $\frac{1}{N_{g}}\sum_{i=1}^{N_{g}}(\frac{1}{T_{i}^{pre}}\sum_{t=1}^{T_{i}^{pre}}X_{it})$, where $i$ denotes a customer belonging to group $g$, $t$ denotes the year-month index, $T_{i}^{pre}$ denotes the length of the consumer $i$’s panel during the pre-adoption period, and $N_{g}$ is the number of customers in group $g$. We next summarize the post-online adoption behaviors in a similar way using the post-online adoption data under each case study. The post-online adoption means for each group ggevent adopters, organic adopters, and offline-only – are calculated as follows: $\frac{1}{N_{g}}\sum_{i=1}^{N_{g}}(\frac{1}{T_{i}^{post}}\sum_{t=1}^{T_{i}^{post}}X_{it})$, where $i$ denotes a customer belonging to group $g$, $t$ denotes the year-month index, $T_{i}^{post}$ denotes the length of the consumer $i$’s panel during the post-adoption period, and $N_{g}$ is the number of customers in group $g$.
B.1 Event 1: COVID-19
In this section, we present the descriptive statistics for customers in the sample for the COVID-19. We first summarize demographics in Table A2. Next, in Tables A3 and A4, we show customer-level summary statistics of average purchase variables over the pre-adoption and post-adoption period, respectively. Finally, in Table A5, we present more detailed summary statistics (e.g., median and quartiles) for all three groups during the post-online-adoption periods (i.e., between May 1st, 2020 and April 30, 2023). Note that the counts of organic adopters and covid adopters from Table A5 slightly differ from the raw count based on our cohorts’ definition because 469 (3.66% of) organic adopters and 933 (2.81% of) covid adopters churned before 2020-05-01, and therefore are not observed in the post-COVID-19 panel. Including these customers does not affect our identification of post-online-adoption ATT, and the estimation results remain the same.
@@ B.1 Event 1: COVID-19
B.2 Event 2: Black Friday
In this section, we present the descriptive statistics for customers in the sample for the Black Friday event. We first summarize demographics in Table A6. Next, in Tables A7 and A8, we show customer-level summary statistics of average purchase variables over the pre-adoption and post-adoption period, respectively. Finally, in Table A9, we present more detailed summary statistics (e.g., median and quartiles) for all three groups during the post-online-adoption periods (i.e., between December 1st, 2019 and October 31, 2020). Note that the counts of organic adopters and Black Friday adopters from Table A9 slightly differ from the raw count based on our cohorts’ definition because 79 (2.58% of) organic adopters and 24 (1.74% of) Black Friday adopters churned before 2019-12-01, and therefore are not observed in the post-Black Friday panel. Including these customers does not affect our identification of post-online-adoption ATT, and the estimation results remain the same.
@@ B.2 Event 2: Black Friday
B.3 Event 3: Loyalty Program
In this section, we present the descriptive statistics for customers in the sample for the loyalty program analysis. We first summarize demographics in Table A10. Next, in Tables A11 and A12, we show customer-level summary statistics of average purchase variables over the pre-adoption and post-adoption period, respectively. Finally, in Table A13, we present more detailed summary statistics (e.g., median and quartiles) for all three groups during the post-online-adoption periods (i.e., between October 1st, 2023 and March 31, 2024)
@@ B.3 Event 3: Loyalty Program
B.4 Loyalty Program Control Variables
@@
    $\%Month\_WithActivation_{it}$: The proportion of months in which customer $i$ has activated the offer up to (but not including) month $t$, calculated as the total number of activation months up to $t$ divided by the number of months since first activation (excluding $t$ itself). For the first month of activation, this variable is defined as 0 (since the denominator is 0); all first-month effects are captured by $isFirstActivation_{it}$. Starting from the second month post-activation, this variable reflects the historical activation rate prior to month $t$.
Appendix C Appendix to the Profitability Analysis: Missing Value Imputation
@@
Appendix D Appendix to Mechanisms, Theoretical Underpinnings and Managerial Implications
D.1 Appendix to the Share of Offline-Online Spend
@@
Table A17: Differential Trends in the Share of Offline Spend (Estimation Results for Equation (A3)) – Event Adopters vs. Organic Adopters
@@
Table A14 presents the DID results of offline spend between adopters and offline-only customers. Table A15 and A16 shows the DID results of offline spend and share of offline spend between event adopters and organic adopters. Table A17 presents the estimation results for the linear trend model by comparing a type of event adopters with organic adopters. Table A18 presents the changes in share of offline spend by different age groups, showing senior adopters tend to have a slower rate of shifting their spend to the online channel.
In §7.2.2, we find that the share of offline spending is persistently higher for covid adopters compared to organic adopters in the post-adoption period, and the downward trend is steeper for organic adopters, indicating that covid adopters shift their spending to online channels slower than organic adopters. To formally test whether these differential trends in share of offline spend are systematic, we estimate the following model with linear time trend on the post-adoption monthly panel data for both groups:
@@
(a) Senior Adopters (Age $\geq 41$)
(b) Young Adopters (Age <41<41)
D.2 Appendix to the Other Metrics
Finally, to further explore the differences in the purchase behavior of different groups of adopters relative to organic adopters post-online-adoption, we estimate the same DID model (as shown in Equation (3) and (4)) on other metrics of interest (Quantity, Number of orders, Unique items, Brands, Categories, and Subcategories). For each study (Covid, Black Friday, and Loyalty Program), we present results for both total and offline metrics. Specifically, total metrics are reported in Tables A19, A21, and A23, while offline metrics appear in Tables A20, A22, and A24, respectively.